# Paying and Persuading

### v4

@@ 2.2 Dynamics
Suppose the stage game is played over many periods, $\xi=0,1,2,\dots$ between Sender and Receiver, both of whom discount the future at a common rate $\delta\in(0,1)$. Moreover, suppose Sender can commit to an entire history-contingent plan of action at the start of the game.
@@
Receiver’s past actions are perfectly observed by Sender, but past states are never observed [^10] Unobservability of past states does not affect the conclusion of Theorem 1 in the i.i.d. case. Relaxing it substantially complicates the remainder of the analysis otherwise, so I maintain it for simplicity.. Thus, a time-$\xi$ history is a sequence $h^{\xi}=\{s_{\zeta},t_{\zeta},a_{\zeta}\}_{\zeta=0}^{\xi-1}$ of past signals, transfers, and actions [^11] I suppose players do not observe their payoffs, so that the state is not revealed at the end of each period. When the state is drawn i.i.d., as in the leading example, this assumption is without loss of generality.. However, because transfers at time $\xi$ are measurable in history $h^{\xi-1}$ and the time-$\xi$ signals and actions and Sender has commitment, it is without loss of generality to set histories to be sequences $\{s_{\zeta},a_{\zeta}\}_{\zeta=0}^{\xi-1}$ of signals and actions only. Let $H^{\xi}=(S\times A)^{\xi}$ be the set of all time-$\xi$ histories and $H=\bigcup_{\xi\in\mathbb{N}}H^{\xi}$ the set of all histories, which we identify with $H^{\infty}=(S\times A)^{\infty}$, the set of infinite histories in the standard way (see Mertens et al. (2015)).
Strategies are tuples $\sigma=(\sigma^{S},\sigma^{R})$ of functions $\sigma^{S}:H\to\Delta(S)^{\Theta}\times\mathcal{T}$ and $\sigma^{R}:H\to\Delta(A)^{S\times\mathcal{T}}$ specifying stage-game strategies for Sender and Receiver at each history. Let $\Sigma=(\Sigma^{S},\Sigma^{R})$ be the set of all strategies. Endow $\Sigma$ with the strong topology (see Appendix B: The Strong Topology on Strategies ). For any strategy $\sigma$, let $\mathbb{P}^{\sigma}\in\Delta(H)$ be the distribution over histories it induces and $\mathbb{Q}^{\sigma}\in(\Delta(\Theta\times A))^{\infty}$ be the joint distribution over outcomes it induces. Let $\mu(h^{\xi};\sigma)$ be Receiver’s time-$\xi$ prior belief about the state at history $h^{\xi}$ given strategy $\sigma$, before any information is disclosed at time $\xi$. In particular, this implies the distribution of beliefs induced by $\sigma^{S}$ at $h^{\xi}$ must be $\mu(h^{\xi};\sigma)$-Bayes plausible.
Let $\sigma(\cdot|h^{\xi})=(\sigma^{S}(\cdot|h^{\xi}),\sigma^{R}(\cdot|h^{\xi}))$ denote the continuation strategies at $h^{\xi}$ and $\sigma(h^{\xi})=(\sigma^{S}(h^{\xi}),\sigma^{R}(h^{\xi}))$ the strategy profile played at $h^{\xi}$. Given some $\sigma$ and $\mathbb{P}^{\sigma}$, say $\sigma$ is obedient if at every $h^{\xi}\in\text{supp}(\mathbb{P}^{\sigma})$, Receiver’s continuation strategy $\sigma^{R}(\cdot|h^{\xi})$ is optimal among the set of all possible continuation strategies:
@@
where $\mathcal{U}(\sigma^{S}(h^{\xi}),\sigma^{R}(h^{\xi}))$ is Receiver’s stage-game payoff given strategies at $h^{\xi}$. A strategy profile $\sigma$ is optimal if it is obedient and attains the maximum of Sender’s payoff among the set of all obedient payoffs. Let $\mathcal{V}^{*}(\mu_{0},\delta)$ be Sender’s payoff at some optimal strategy when the prior is $\mu_{0}$ and players discount the future at $\delta$.
@@
Let $\sigma$ be obedient. There exists obedient $\sigma^{*}$ with $S=\Delta(\Theta)\times A$ such that for each signal $s$ reached on the path of play, $\mathbb{E}_{\theta|(\mu,a)}=\mu$, $\sigma^{R}(h^{\xi})(\mu,a)=a$, and $t((\mu,a),a^{\prime})\neq 0\implies a=a^{\prime}$ and moreover $\mathbb{P}^{\sigma}=\mathbb{P}^{\sigma^{*}}.$
The proof is straightforward but deferred to PROOF OF LEMMA 2 Lemma 2 implies that for each obedient strategy there is an outcome equivalent (and hence payoff equivalent) strategy $\sigma^{*}$ which is direct in beliefs and actions: it tells Receiver (directly) what belief to hold, and then recommends an incentive compatible action. Moreover, it further simplifies the space of transfers: Sender need only pay Receiver if Receiver follows the recommended strategy. This implies the set of transfers needed to be specified simplifies from one for each pair of actions only to one for each action recommendation, a much lower dimensional space.
@@
Note Lemma 2 implies it is equivalent to recast the problem into finding Bayes plausible distributions of beliefs $\rho\in\Delta(\Delta(\Theta))$, transfer rules $t:S\to\mathbb{R}_{+}$, and recommended (random) actions $\alpha:\Delta(\Theta)\to\Delta(A)$ such that (1) Receiver finds it incentive compatible to take all actions $a\in\text{supp}(\alpha(\mu))$, and (2) beliefs are Bayes plausible, $\mathbb{E}_{\rho}[\mu]=\mu_{0}$. Paired with a future utility continuation promise $u^{\prime}:A\times\Delta(\Theta)\to\mathbb{R}$, this implies the following recursive formulation of the optimal strategy:
$V(\bar{u},\mu_{0},\delta)=\max_{\{\rho,t,u^{\prime},\alpha\}}\Big[\mathbb{E}_{\rho}\Big[(1-\delta)\,\mathbb{E}_{\mu,\alpha}[v(a,\theta)$ $-kt(\mu,a)]+\delta V(u^{\prime}(\mu,a),M\mu,\delta)\Big]\Big]$ (FE)
@@
$\mathbb{E}_{\rho}[(1-\delta)\,\mathbb{E}_{\mu,\alpha}[u(a,\theta)+t(\mu,a)]+\delta u^{\prime}(\mu,a)]$ $\geq\bar{u}$ (PK)
@@
$\max\{(1-\delta)t(\mu,a),\,u^{\prime}(\mu,a)\}\in[-C,C]\quad\forall$ $\mu\in\text{supp}(\rho),a\in\text{supp}(\alpha(\mu)),$ (BD)
@@
Proposition 1 formalizes the way in which solving for the functional equation above implies finding an optimal $\sigma$. Throughout, represent $\sigma^{S}(h^{\xi})$ as a tuple $\{\rho(h^{\xi}),t(h^{\xi})\}$ and Receiver action $\sigma^{R}(h^{\xi})$ as a random action $\{\alpha(h^{\xi})\}$, and let $\mathcal{U}(h^{\xi};\sigma)$ be Receiver’s continuation utility at history $h^{\xi}$. When there is any ambiguity about which strategy $\{\rho(\cdot),t(\cdot),\alpha(\cdot)\}$ represents, I will add an additional argument to specify explicitly, i.e. $\{\rho(\cdot;\sigma),t(\cdot;\sigma),\alpha(\cdot;\sigma)\}$ though I suppress the extra notation in the absence of such ambiguity.
@@
(FE) gives the value of Sender’s problem: $V(0,\mu_{0},\delta)=\mathcal{V}^{*}(\mu_{0},\delta)$. Moreover, there exists an optimal $\sigma$ where for any $h^{\xi}\in\text{supp}(\rho)$,
@@
Proposition 1 follows from basic principles in dynamic programming and is proven in PROOF OF PROPOSITION 1 Importantly, it implies that it is sufficient to characterize (1) the evolution of solutions to the functional equation at different prior beliefs and utility promises and (2) the evolution of the path of utility promises at the optimum in order to obtain a full characterization of the optimal strategy $\sigma^{*}$. Towards this end, let $\mathcal{F}(u,\mu)$ be the solution to the value function (for an implicit fixed discount rate) at utility promise $u$ and belief $\mu$. I conclude the description of the model by stating some basic facts about the value function.
@@
There exists a unique $V$ solving (FE), and moreover $V$ is continuous in $u$ over [C,C][-C,C] and $\mu_{0}$ on $\Delta(\Theta)$. Finally, $V$ is concave nondecreasing in $u$, and its right derivative satisfies $V_{+}^{\prime}(u,\mu)\geq-k$ for any $u,\mu$.
@@ 3 Time Structure of Payments
Transfers are a last resort at $\sigma$ if, for any $h^{\xi}\in\mathbb{P}^{\sigma}$, $t(h^{\xi})(\mu,a)>0$ implies $\rho(h^{\zeta})$ reveals the state at every on-path successor history $h^{\zeta}\succsim\{h^{\xi},\mu,a\}$.
@@
There is an optimal strategy $\sigma^{*}$ where transfers are a last resort.
@@
Theorem 1 is the main takeaway of this paper. It gives a characterization of when payments should be used to dynamically motivate Receiver in the prescence of informational incentives. If Sender ever wants to motivate an agent via transfers, they should instead turn to dynamic information first. Only after Sender has completely “drawn down” on their stock of information (by promising full information at all continuation histories) should they start using transfers to motivate the agent.
@@
The intuition behind the result is as follows. Suppose Sender at any point wishes to motivate the agent to take an action (fixing an induced belief today). They can do this in two ways: by paying the agent, or by giving them more information. Since more information makes Receiver’s decision more efficient, Sender can always give a little more information to increase Receiver’s utility (and possibly Sender’s at some beliefs). From here, at any of the newly induced posterior beliefs where Receiver’s new action adversely affects Sender’s utility, they can pay Receiver to take an action equal to their original action. The amount they have to pay is exactly equal to the change in Receiver’s payoff along this new posterior relative to the new action, and thus comes at a efficiency cost of at most rate k-k. Such a future joint pay-and-persuade scheme cannot be worse off than simply paying the agent today (and will do strictly better so long as more information benefits Sender too), and hence so long as such a scheme is feasible at some future continuation history, can be chosen in lieu of payments today along the optimal path. This observation is the substance of the pullback and squeezing lemmas (Lemmas 4 and 5). The remainder of the proof handles the technicalities that arise given the generality of the model and the infinite set of histories.
@@
It is worth explicitly flagging that the above informal argument is distinct from two alternative intuitions one might have about the result. First is the intuition that transfers must be backloaded because information provision is “always free” (since there are no costs to information acquisition). In fact, it can be that in the absence of transfers, the cost of giving the agent more information can be arbitrarily high because it allows Receiver to take actions which can be very damaging to Sender but which only marginally benefit Receiver. With transfers, however, the pullback lemma shows that Sender can bound the cost of transferring utility to Receiver via information and transfers by at most k-k, which is the marginal cost of simply paying an agent to fulfill a utility promise. Second is the intuition that transfers can be mechanically backloaded because payments are perfectly intertemporally substitutable. While it is true payments can be mechanically pushed backwards (at the cost of increasing them by $\frac{1}{\delta}$) while maintaining optimality, this mechanical push-back does not change the information structure. Thus, there is no guarantee after pushing back payments that (1) for all optimal contracts the state is fully revealed or that (2) pushing transfers back would encourage information revelation. Thus, there is no way to guarantee by simply backloading transfers that they would ever be used as a last resort in some optimal contract.
@@
Finally, note Theorem 1 does not guarantee every optimal contract strictly backloads transfers; only that one such contract exists. In Section 4.3 (see Proposition 6), I expand on this indeterminancy and give conditions under which transfers are strictly backloaded—there are always periods where dynamic information design is strictly preferred to transfers as a way of motivating Receiver.
@@
The second step: once we are at the steepest part of the Pareto frontier, it is without loss of optimality to provide incentives by giving Receiver full information with some probability. To state this step we need a bit more notation. At any time $\xi$, let $\mathbb{Q}_{\xi}^{\sigma}\in\Delta(\Theta\times A)$ be the unconditional probability distribution over outcomes (averaging over histories) induced by strategy $\sigma$ at time $\xi$, with $\mathbb{Q}_{\xi}^{\sigma}(\theta)\in\Delta(A)$ the distribution over actions fixing some state. Moreover, define
$u^{RFI}(\mu)=(1-\delta)\sum_{\xi=0}^{\infty}\delta^{\xi}\mathbb{E}_{M^{\xi}\mu}[\max_{a\in A}u(a,\theta)]$
to be Receiver’s expected full-information payoff, starting at belief $\mu$. Finally, say $\sigma$ is $u$-constrained optimal if it is optimal among all obedient $\sigma$ which guarantee Receiver a payoff of at least $u$. Note all continuation strategies $\sigma(\cdot|h^{\xi})$ are $\mathcal{U}(h^{\xi}|\sigma)$-constrained optimal. We can now state the following lemma, proven in the PROOF OF LEMMA 4
@@
Fix $(u,\mu_{0})$ such that $V_{+}^{\prime}(u,\mu_{0})=-k$, and let $\sigma$ be $u$-constrained optimal. Then there exists obedient $\sigma^{*}$ such that
@@
The pullback lemma combined with Lemma 3 implies that at any history where transfers are used, it is possible to further increase utility by simply revealing the state with positive probability. This comes at a cost of $k$, and hence is an efficient way to transfer utility once $V_{+}^{\prime}(u,\mu_{0})=-k$. This implies that full information is on the Pareto frontier at some point, since it is always efficient to provide utility by giving a little bit more information (i.e. an increase in $\tilde{u}$ corresponds to an increase in $\alpha$).
@@
where $\rho^{FI}$ is the experiment that reveals the state.
@@
The proof can be found in PROOF OF LEMMA 5 The pullback and squeezing lemmas give an iterative procedure we can use to obtain an optimal strategy where transfers are a last resort. Start with some optimum $\sigma_{0}$. Define the function
@@
    For some finite $N$, $\varphi(\sigma_{N})=\infty$. Then $\sigma_{N}$ is an optimal strategy where transfers are a last resort, and we are done.
@@
In the latter case, a compactness argument gives a convergent subsequence $\{\sigma_{k_{n}}\}\subset\{\sigma_{k}\}$ in the strong topology. Let $\sigma_{\infty}$ be this subsequential limit. Since optimal strategies are closed (because $\mathbb{P}^{\sigma}$ is continuous in $\sigma$ by Appendix B: The Strong Topology on Strategies ), $\sigma_{\infty}$ is also optimal. But $\varphi(\sigma_{\infty})=\infty$ must be true, and so $\sigma_{\infty}$ is an optimal strategy where transfers are a last resort. ∎

### v5

@@ 2.2 Dynamics
Suppose the stage game is played over many periods, $\xi=0,1,2,\dots$ between Sender and Receiver, both of whom discount the future at a common rate $\delta\in(0,1)$. Moreover, suppose Sender can commit to an entire history-contingent plan of action at the start of the game. Call this plan Sender’s contract.
@@
Receiver’s past actions are perfectly observed by Sender, but past states are never observed. [^12] Unobservability of past states does not affect the conclusion of Theorem 1 in the i.i.d. case. Relaxing it substantially complicates the remainder of the analysis otherwise, so I maintain it for simplicity. Thus, a time-$\xi$ history is a sequence $h^{\xi}=\{s_{\zeta},t_{\zeta},a_{\zeta}\}_{\zeta=0}^{\xi-1}$ of past signals, transfers, and actions. [^13] I suppose players do not observe their payoffs, so that the state is not revealed at the end of each period. When the state is drawn i.i.d., as in the leading example, this assumption is without loss of generality. However, because transfers at time $\xi$ are measurable in history $h^{\xi-1}$ and the time-$\xi$ signals and actions and Sender has commitment, it is without loss of generality to set histories to be sequences $\{s_{\zeta},a_{\zeta}\}_{\zeta=0}^{\xi-1}$ of signals and actions only. Let $H^{\xi}=(S\times A)^{\xi}$ be the set of all time-$\xi$ histories and $H=\bigcup_{\xi\in\mathbb{N}}H^{\xi}$ the set of all histories, which I identify with $H^{\infty}=(S\times A)^{\infty}$, the set of infinite histories in the standard way (see Mertens et al. (2015)).
A contract is a tuple $\sigma=(\sigma^{S},\sigma^{R})$ of functions $\sigma^{S}:H\to\Delta(S)^{\Theta}\times\mathcal{T}$ and $\sigma^{R}:H\to\Delta(A)^{S\times\mathcal{T}}$ specifying stage-game strategies for Sender and Receiver at each history. Let $\Sigma=(\Sigma^{S},\Sigma^{R})$ be the set of all contracts. Endow $\Sigma$ with the strong topology (see Appendix B: The Strong Topology on Contracts ). For any contract $\sigma$, let $\mathbb{P}^{\sigma}\in\Delta(H)$ be the distribution over histories it induces and $\mathbb{Q}^{\sigma}\in(\Delta(\Theta\times A))^{\infty}$ be the joint distribution over outcomes it induces. Let $\mu(h^{\xi};\sigma)$ be Receiver’s time-$\xi$ prior belief about the state at history $h^{\xi}$ given contract $\sigma$, before any information is disclosed at time $\xi$. In particular, this implies the distribution of beliefs induced by $\sigma^{S}$ at $h^{\xi}$ must be $\mu(h^{\xi};\sigma)$-Bayes plausible.
Let $\sigma(\cdot|h^{\xi})=(\sigma^{S}(\cdot|h^{\xi}),\sigma^{R}(\cdot|h^{\xi}))$ denote the continuation contracts at $h^{\xi}$ and $\sigma(h^{\xi})=(\sigma^{S}(h^{\xi}),\sigma^{R}(h^{\xi}))$ the profile played at $h^{\xi}$. Given some $\sigma$ and $\mathbb{P}^{\sigma}$, say $\sigma$ is obedient if at every $h^{\xi}\in\text{supp}(\mathbb{P}^{\sigma})$, Receiver’s continuation contract $\sigma^{R}(\cdot|h^{\xi})$ is optimal among the set of all possible continuation contracts:
@@
where $\mathcal{U}(\sigma^{S}(\cdot|h^{\xi}),\sigma^{R}(\cdot|h^{\xi}))$ is Receiver’s continuation payoff given contracts at $h^{\xi}$. Note it is without loss of generality to consider contracts where off-path deviations are punished with no information and zero transfers forevermore. A contract $\sigma$ is optimal if it is obedient and attains the maximum of Sender’s payoff among all obedient contracts. Let $\mathcal{V}^{*}(\mu_{0},\delta)$ be Sender’s payoff at optimum when the prior is $\mu_{0}$ and players discount the future at $\delta$.
@@
Let $\sigma$ be obedient. There exists obedient $\sigma^{*}$ with $S=\Delta(\Theta)\times A$ such that for all $h^{\xi}\in\text{supp}(\mathbb{P}^{\sigma^{*}})$, $\mu(h^{\xi},\sigma^{*}|(\mu,a))=\mu$, $\sigma^{*R}(h^{\xi})(\mu,a)=a$, and $t((\mu,a),a^{\prime};\sigma^{*})\neq 0\implies a=a^{\prime}$ such that $\mathbb{Q}^{\sigma}=\mathbb{Q}^{\sigma^{*}}$.
The proof is a standard relabeling argument, done in PROOF OF LEMMA 2 Lemma 2 implies that for each obedient contract there is an outcome equivalent (and hence payoff equivalent) contract $\sigma^{*}$ which is direct in beliefs and actions: it tells Receiver (directly) what belief to hold, and then recommends an incentive compatible action. Moreover, it further simplifies the space of transfers: Sender need only pay Receiver if Receiver follows the recommended action. This implies the set of transfers needed to be specified simplifies from one for each pair of actions only to one for each action recommendation, a much lower dimensional space.
@@
Note Lemma 2 implies it is equivalent to recast the problem into finding Bayes plausible distributions of beliefs $\rho\in\Delta(\Delta(\Theta))$, transfer rules $t:\Delta(\Theta)\times A\to\mathbb{R}_{+}$, and recommended (random) actions $\alpha:\Delta(\Theta)\to\Delta(A)$ such that (1) Receiver finds it incentive compatible to take all actions $a\in\text{supp}(\alpha(\mu))$, and (2) beliefs are Bayes plausible, $\mathbb{E}_{\rho}[\mu]=\mu_{0}$. Paired with a future utility continuation promise $u^{\prime}:\Delta(\Theta)\times A\to\mathbb{R}$, this implies the following recursive formulation of the optimal contract:
$V(\bar{u},\mu_{0},\delta)=\max_{\{\rho,t,u^{\prime},\alpha\}}\Big\{\mathbb{E}_{\rho}\Big[\mathbb{E}_{\mu,\alpha}[(1-\delta)\,[v(a,\theta)$ $-kt(\mu,a)]+\delta V(u^{\prime}(\mu,a),M\mu,\delta)]\Big]\Big\}$ (FE)
@@
$\mathbb{E}_{\rho}[\mathbb{E}_{\mu,\alpha}[(1-\delta)[u(a,\theta)+t(\mu,a)]+\delta u^{\prime}(\mu,a)]]$ $\geq\bar{u}$ (PK)
@@
$\max\{(1-\delta)t(\mu,a),\,|u^{\prime}(\mu,a)|\}\in[0,C]\quad\forall$ $\mu\in\text{supp}(\rho),a\in\text{supp}(\alpha(\mu)),$ (BD)
@@
Proposition 1 formalizes the way in which solving for the functional equation above implies finding an optimal $\sigma$. Throughout, represent $\sigma^{S}(h^{\xi})$ as a tuple $\{\rho(h^{\xi}),t(h^{\xi})\}$ and Receiver action $\sigma^{R}(h^{\xi})$ as a random action $\{\alpha(h^{\xi})\}$, and let $\mathcal{U}((h^{\xi},\mu,a);\sigma)$ be Receiver’s continuation utility at history $h^{\xi}$ after forming posterior belief $\mu$ and taking action $a$ (both assumed to be on the path of play). When there is any ambiguity about which contract $\{\rho(\cdot),t(\cdot),\alpha(\cdot)\}$ represents, I will add an additional argument to specify explicitly, i.e. $\{\rho(\cdot;\sigma),t(\cdot;\sigma),\alpha(\cdot;\sigma)\}$ though I suppress the extra notation in the absence of such ambiguity.
@@
(FE) gives the value of Sender’s problem: $V(0,\mu_{0},\delta)=\mathcal{V}^{*}(\mu_{0},\delta)$. Moreover, there exists an optimal $\sigma$ where for any $h^{\xi}\in\text{supp}(\mathbb{P}^{\sigma})$,
@@
Proposition 1 follows from basic principles in dynamic programming and is proven in PROOF OF PROPOSITION 1 Importantly, it implies that it is sufficient to characterize (1) the evolution of the argument maximizers to the functional equation at different prior beliefs and utility promises and (2) the evolution of the path of utility promises at the optimum in order to obtain a full characterization of the optimal contract $\sigma^{*}$. Towards this end, let $\mathcal{F}(u,\mu)$ be the collection of all tuples which solve the value function program (for an implicit fixed discount rate) at utility promise $u$ and belief $\mu$. I conclude the description of the model by stating some basic facts about the value function.
@@
There exists a unique $V$ solving (FE), and moreover $V$ is continuous in $u$ over [C,C][-C,C] and $\mu_{0}$ on $\Delta(\Theta)$. Finally, $V$ is concave nonincreasing in $u$, and its right derivative satisfies $V_{+}^{\prime}(u,\mu)\geq-k$ for any $u,\mu$.
@@ 3 Time Structure of Payments
Transfers are a last resort at $\sigma$ if, for any $h^{\xi}\in\text{supp}(\mathbb{P}^{\sigma})$, $t(h^{\xi})(\mu,a)>0$ implies $\rho(h^{\zeta})$ reveals the state at every on-path successor history $h^{\zeta}\succsim\{h^{\xi},\mu,a\}$.
@@
There is an optimal contract $\sigma^{*}$ where transfers are a last resort.
@@
Theorem 1 is the main takeaway of this paper. It gives a characterization of when payments should be used to dynamically motivate Receiver in the presence of informational incentives. If Sender ever wants to motivate an agent via transfers, they should instead turn to dynamic information first. Along the optimal contract where transfers are a last resort, Sender should start using transfers to motivate the agent only after they have completely “drawn down” on their stock of information (by promising full information at all continuation histories).
@@
The intuition behind the result is as follows. Suppose Sender at any point wishes to motivate the agent to take an action (fixing an induced belief today). They can do this in two ways: by paying the agent, or by promising them more information in the future. Since more information makes Receiver’s decision more efficient, Sender can always give a little more information to increase Receiver’s utility (and possibly Sender’s at some beliefs). From here, at any of the newly induced posterior beliefs where Receiver’s new action adversely affects Sender’s utility, they can pay Receiver to take an action equal to their original action. The amount they have to pay is exactly equal to the change in Receiver’s payoff along this new posterior relative to the new action, and thus comes at a efficiency cost of at most rate $k$. Such a future joint pay-and-persuade scheme cannot be worse off than simply paying the agent today (and will do strictly better so long as more information benefits Sender too), and hence so long as such a scheme is feasible at some future continuation history, can be chosen in lieu of payments today along the optimal path. This observation is the substance of the pullback and squeezing lemmas (Lemmas 4 and 5). The remainder of the proof handles the technicalities that arise given the generality of the model and the infinite set of histories.
@@
It is worth explicitly flagging that the above informal argument is distinct from two alternative intuitions one might have about the result. First is the intuition that transfers must be backloaded because information provision is “always free” (since there are no costs to information acquisition). In fact, it can be that in the absence of transfers, the cost of giving the agent more information can be arbitrarily high because it allows Receiver to take actions which can be very damaging to Sender but which only marginally benefit Receiver. With transfers, however, the pullback lemma shows that Sender can bound the cost of transferring utility to Receiver via information and transfers by at most $k$, which is the marginal cost of simply paying an agent to fulfill a utility promise. Second is the intuition that transfers can be mechanically backloaded because payments are perfectly intertemporally substitutable. While it is true payments can be mechanically pushed backwards (at the cost of increasing them by $\frac{1}{\delta}$) while maintaining optimality, this mechanical push-back does not change the information structure. Thus, there is no guarantee after pushing back payments that (1) for all optimal contracts the state is fully revealed or that (2) pushing transfers back would encourage information revelation. Thus, there is no way to guarantee by simply backloading transfers that they would ever be used as a last resort in some optimal contract. In contrast, my construction (in Lemma 5) (1) pushes transfers into specific future histories where the state is not fully revealed, and (2) simultaneously reveals more information to increase the agents’ indirect consumption utility. Combined, this alleviates the need for transfers today, while also guaranteeing that future histories see more informative experiments. Finally, transfers are guaranteed to be a last resort by appealing to an appropriate compactness argument, after having pushed utility far enough back that all future experiments fully reveal the state (Lemma 5).
@@
Finally, note Theorem 1 does not guarantee every optimal contract strictly backloads transfers; only that one such contract exists. In Sections 4.3 and 4.4 (see Propositions 6 and 8), I expand on this indeterminancy and give conditions under which transfers are strictly backloaded—there are always periods where dynamic information design is strictly preferred to transfers as a way of motivating Receiver.
@@
The second step: once we are at the steepest part of the Pareto frontier, it is without loss of optimality to provide incentives by giving Receiver full information with some probability. To state this step we need a bit more notation. At any time $\xi$, let $\mathbb{Q}_{\xi}^{\sigma}\in\Delta(\Theta\times A)$ be the unconditional probability distribution over outcomes (averaging over histories) induced by contract $\sigma$ at time $\xi$, with $\mathbb{Q}_{\xi}^{\sigma}(\theta)\in\Delta(A)$ the distribution over actions fixing some state. Moreover, define
$u^{RFI}(\mu)=(1-\delta)\sum_{\xi=0}^{\infty}\delta^{\xi}\mathbb{E}_{M^{\xi}\mu}\left[\max_{a\in A}u(a,\theta)\right]$
to be Receiver’s expected full-information payoff, starting at belief $\mu$. Finally, say $\sigma$ is $u$-constrained optimal if it is optimal among all obedient $\sigma$ which guarantee Receiver a payoff of at least $u$. Let $mu(h^{\xi},\sigma)$ be the posterior belief over states given history $h^{\xi}$ and strategy $\sigma$. Note all continuation contracts $\sigma(\cdot|h^{\xi})$ are $\mathcal{U}(h^{\xi}|\sigma)$-constrained optimal. We can now state the following lemma, proven in the PROOF OF LEMMA 4
@@
Fix $(u,\mu_{0})=(\mathcal{U}(h^{\xi}|\sigma),\mu(h^{\xi},\sigma))$ such that $V_{+}^{\prime}(u,\mu_{0})=-k$, and let $\sigma$ be $u$-constrained optimal. Then there exists obedient $\sigma^{*}$ such that
@@
The pullback lemma combined with Lemma 3 implies that at any history where transfers are used, it is possible to further increase utility by simply revealing the state with positive probability. This comes at a cost of (at most) $k$, and hence is an efficient way to transfer utility once $V_{+}^{\prime}(u,\mu_{0})=-k$. This implies that full information is on the Pareto frontier at some point, since it is always efficient to provide utility by giving a little bit more information (i.e. an increase in $\tilde{u}$ corresponds to an increase in $\alpha$).
@@
where $\rho^{FI}$ is the experiment that always reveals the state.
@@
The proof can be found in PROOF OF LEMMA 5 The pullback and squeezing lemmas give an iterative procedure we can use to obtain an optimal contract where transfers are a last resort. Start with some optimum $\sigma_{0}$. Define the function
@@
    For some finite $N$, $\varphi(\sigma_{N})=\infty$. Then $\sigma_{N}$ is an optimal contract where transfers are a last resort, and we are done.
@@
In the latter case, a compactness argument gives a convergent subsequence $\{\sigma_{k_{n}}\}\subset\{\sigma_{k}\}$ in the strong topology. Let $\sigma_{\infty}$ be this subsequential limit. Since optimal contracts are closed (because $\mathbb{P}^{\sigma}$ is continuous in $\sigma$ by Appendix B: The Strong Topology on Contracts ), $\sigma_{\infty}$ is also optimal. But $\varphi(\sigma_{\infty})=\infty$ must be true, and so $\sigma_{\infty}$ is an optimal contract where transfers are a last resort. ∎